OpenStudy (peach3113):

Stanley has 200 chocolate bars. He eats 8/9. What does Stanley have now?

OpenStudy (peach3113):

Nope.

OpenStudy (peach3113):

I know enough about math to know that is incorrect.

OpenStudy (anonymous):

Divide by 8?

OpenStudy (anonymous):

divide 9 and 8

OpenStudy (peach3113):

I think it's 200/1 divided by 8/9

OpenStudy (anonymous):

then divide 200 by the answer

OpenStudy (peach3113):

I just want to know if that's right.

OpenStudy (anonymous):

^That would make more sense XD @Peach3113

OpenStudy (anonymous):

As Yeah it correct

OpenStudy (anonymous):

8/9 is a fraction, and 200 is the whole. You need to divide it and turn 200 into a fraction so 200/1 / 8/9. Is that wrong?

OpenStudy (peach3113):

I think that'd be right, thanks AnimeWover.

OpenStudy (mathstudent55):

You are trying to find 8/9 of 200. In math, "of" usually means "times." 8/9 of 200 means 8/9 times 200. \(\dfrac{8}{9} \times 200 = \dfrac{8}{9} \times \dfrac{200}{1} = \dfrac{8 \times 200}{9 \times 1} \) Can you finish now?

OpenStudy (anonymous):

I am actually surprised I did this one. XD. I am horrible at fractions

OpenStudy (peach3113):

Okay, yes, thank you @mathstudent55

OpenStudy (anonymous):

I'm wrong?

OpenStudy (anonymous):

Figured

OpenStudy (mathstudent55):

@AnimeWover You are dividing. The problem needs multiplication.

OpenStudy (peach3113):

800/1= 800. How is that correct?

OpenStudy (peach3113):

He only had 200.

OpenStudy (anonymous):

I understand that diving fractions is basically multiplying. But I also got 800/1 = 800. Um?

OpenStudy (peach3113):

@AnimeWover Yeah idk. XD I'm losing my mind over this question.

OpenStudy (anonymous):

wait 200 x 8 isn't 800

OpenStudy (anonymous):

200 x 8 = 1000.

OpenStudy (peach3113):

OH

OpenStudy (anonymous):

XD

OpenStudy (peach3113):

No, 1600

OpenStudy (peach3113):

I'm an idiot. XD

OpenStudy (anonymous):

OH FAIL

OpenStudy (anonymous):

I counted to 5 not 8 XD

OpenStudy (anonymous):

1600 / 9???

OpenStudy (anonymous):

177 is 1600 / 9....

OpenStudy (peach3113):

177.something

OpenStudy (mathstudent55):

Here is the original question: "Stanley has 200 chocolate bars. He eats 8/9. What does Stanley have now?" Here is what it means: "Stanley has 200 chocolate bars. He eats 8/9 of the 200 chocolate bars. What does Stanley have now?" The answer is "Now he has 1/9 of the 200 chocolate bars." You can use the calculation I did above to find 8/9 of the bars, and then subtract that number form 200, or you can just calculate 1/9 of 200.

OpenStudy (peach3113):

I just know that stanley must have diabetes

OpenStudy (anonymous):

LOL

OpenStudy (peach3113):

That is a lot of chocolate.

OpenStudy (anonymous):

200/9 = the answer It is that simple

OpenStudy (peach3113):

I'll figure this out eventually. And hopefully so will stanley.

OpenStudy (anonymous):

mayve 8/9 bars from a hershey sized bar?

OpenStudy (peach3113):

He needs help.

OpenStudy (anonymous):

200 / 9 = 22. Stanley has 22 left.

OpenStudy (anonymous):

Um...I think i broke the rules.

OpenStudy (peach3113):

Hmm. Thank you Anime. :P

OpenStudy (peach3113):

Oh no, it's fine.

OpenStudy (anonymous):

XD

OpenStudy (peach3113):

I mean, we're both trying to figure this out.

OpenStudy (anonymous):

Omg Stanley must have the illness 0-0

OpenStudy (peach3113):

Tapeworm maybe?

OpenStudy (anonymous):

@Peach3113 can you help me when your finished?

OpenStudy (anonymous):

like omg. Who would eat 178 BARS OF CHOCOLATE?

OpenStudy (anonymous):

@Peach3113 She is helping me :c

OpenStudy (peach3113):

YEAH XD

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

MY READING XD

OpenStudy (peach3113):

XD

OpenStudy (anonymous):

(Pleads with Senpai Peachy)

OpenStudy (peach3113):

It's kay bby.

OpenStudy (peach3113):

I'll halp u

OpenStudy (anonymous):

thank you